### v7

# PAD: Personalized Alignment of LLMs at Decoding-Time

Abstract
Aligning with personalized preferences, which vary significantly across cultural, educational, and political differences, poses a significant challenge due to the computational costs and data demands of traditional alignment methods. In response, this paper presents Personalized Alignment at Decoding-time (PAD), a novel framework designed to align LLM outputs with diverse personalized preferences during the inference phase, eliminating the need for additional training. By introducing a unique personalized reward modeling strategy, this framework decouples the text generation process from personalized preferences, facilitating the generation of generalizable token-level personalized rewards. The PAD algorithm leverages these rewards to guide the decoding process, dynamically tailoring the base model’s predictions to personalized preferences. Extensive experimental results demonstrate that PAD not only outperforms existing training-based alignment methods in terms of aligning with diverse preferences but also shows significant generalizability to preferences unseen during training and scalability across different base models. This work advances the capability of LLMs to meet user needs in real-time applications, presenting a substantial step forward in personalized LLM alignment. Our model and code are available here.
1 Introduction
Recent advancements have demonstrated success in aligning language models with human preferences and values (Stiennon et al., 2020; Bai et al., 2022; Ouyang et al., 2022; Achiam et al., 2023). Representative methods such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) and DPO (Rafailov et al., 2024b) typically optimize a policy model with training signals from an explicit or implicit reward model. The reward model captures the ‘general’ human preferences or values from human feedback. However, in this pluralistic world, users’ preferences can diverge significantly based on their different cultures, educational backgrounds, religions, and political stands (Gordon et al., 2022; Sorensen et al., 2024a; Jang et al., 2023; Cheng et al., 2023). Furthermore, even for the same person, the preference of a particular LLM response can vary when the application scenario changes. Hence, there always exists a proportion of human preferences that cannot be unified by the general preference, also known as personalized preferences, which current alignment frameworks struggle to align with due to the need for high-quality datasets and substantial computational costs in policy optimization.
How can we align with personalized preferences without the need for additional data collection and policy training? In this paper, we introduce Personalized Alignment at Decoding-time (PAD), which aims to align LLM outputs with diverse personalized preferences during the inference phase without requiring additional training. To achieve this, we first propose a personalized reward modeling strategy, which decouples the text generation process (modeled as a Markov Decision Process) from personalized preferences, thereby enabling the acquisition of generalizable token-level personalized rewards. Based on this, we then formulate a personalized reward model (PersRM). During decoding, the PersRM scores the base model’s top-K predictions at each token generation step based on the current generation and personalized preferences. Finally, this score is combined with standard decoding likelihoods to adjust the base model’s predictions. The advantages of PAD are as follows: (1) It requires only a single policy model (i.e., the base model) aligned with general preferences (General Policy), eliminating the need for training additional policy models (Training-free). (2) It utilizes only a single reward model (Single Reward). (3) It does not require pre-defined personalized preferences to generalize to preferences not seen during the training phase (Generalizability). A checklist of PAD’s advantages over previous methods is presented in Table 1.
Table 1: A checklist for key characteristics of previous methods and PAD. “-”: Not Applicable.
Method Training-free General Policy Single Reward Generalizability
MORLHF (Li et al., 2020) $\times$ $\checkmark$ $\times$ $\times$
MODPO (Zhou et al., 2023) $\times$ $\checkmark$ - $\times$
Personalized soups (Jang et al., 2023) $\times$ $\times$ $\times$ $\times$
Preference Prompting (Jang et al., 2023) $\checkmark$ $\checkmark$ - $\checkmark$
Rewarded soups (Rame et al., 2024) $\times$ $\times$ $\times$ $\times$
RiC (Yang et al., 2024b) $\times$ $\checkmark$ $\times$ $\times$
DPA (Wang et al., 2024a) $\times$ $\checkmark$ $\checkmark$ $\times$
Args (Khanov et al., 2024) $\checkmark$ $\checkmark$ $\times$ $\times$
MOD (Shi et al., 2024) $\checkmark$ $\times$ $\times$ $\times$
MetaAligner (Yang et al., 2024a) $\checkmark$ $\checkmark$ - $\checkmark$
PAD (Ours) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Our contributions can be summarized as follows:
  • We propose a novel personalized reward modeling strategy that decouples the dynamics of text generation from personalized preferences. This strategy enables the acquisition of generalizable token-level personalized rewards with a single personalized reward model (PersRM).
  • We propose a novel personalized alignment at decoding-time (PAD) algorithm that performs guided decoding with the guidance of token-level personalized rewards, while not requiring training additional policy models.
  • Extensive experiments demonstrate that PAD outperforms existing training-based methods in aligning with diverse personalized preferences. Furthermore, the results highlight PAD’s effectiveness in generalizing to unseen preferences and its model-agnostic scalability.
2 Related Works
Large language model alignment.
Large language model alignment aims to align LLMs with human preferences. A common approach involves using an RLHF (Reinforcement Learning with Human Feedback) framework Christiano et al. (2017); Bai et al. (2022) where a reward model is trained based on human feedback, and Proximal Policy Optimization (PPO) (Schulman et al., 2017) is employed to derive the aligned policy model. Notably, decoding-time alignment offers an alignment paradigm that does not require expensive RL training (Mudgal et al., 2023; Khanov et al., 2024; Han et al., 2024; Liu et al., 2024; Huang et al., 2024). Controlled Decoding (CD) (Mudgal et al., 2023) utilizes a prefix scorer module trained to assess value functions for rewards, allowing controlled generation from a frozen base model. ARGS (Khanov et al., 2024) proposed using a reward signal to adjust probabilistic predictions, thereby generating semantically aligned texts. DeAL (Huang et al., 2024) focusing on heuristic-guided searches to better meet diverse alignment objectives.
Personalized alignment.
As users exhibit diverse preferences and values for a single task, it is essential to align LLMs to personalized preferences (Kirk et al., 2023; Sorensen et al., 2023; 2024b; Yao et al., 2023; Kirk et al., 2024; Zhong et al., 2024; Han et al., 2024). One line of work achieves joint optimization for different personalized preferences by defining a reward function with multiple dimensions and performing policy optimization (Zhou et al., 2023; Wang et al., 2024a; b; Guo et al., 2024; Yang et al., 2024b; Chakraborty et al., 2024; Sun et al., 2024; Li et al., 2024). Additionally, some approaches merge model parameters or predictions for each dimension to handle their diverse combinations (Jang et al., 2023; Rame et al., 2024; Park et al., 2024; Shi et al., 2024). Lastly, prompt-based methods align personalized preferences by designing diverse prompts or post-processing techniques (Yang et al., 2024a; Lee et al., 2024; Hwang et al., 2023; Jafari et al., 2024).
3 Method
3.1 Preliminaries
In this section, we first define the per-token Markov Decision Process (MDP) for large language models (LLMs) and describe its relationship to classic Reinforcement Learning from Human Feedback (RLHF) approaches. Then, we describe the characteristics and challenges of personalized alignment.
Text generation as token-level markov decision process.
The standard text generation process of large language models (LLMs) with prompt $\mathbf{x}$ and response $\mathbf{y}$ can be defined as a token-level Markov Decision Process (MDP). MDP is denoted as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},R,T)$, where the state space $\mathcal{S}$ consists of the prompt and all tokens generated so far (i.e., $\mathbf{s}_{t}=(\mathbf{x},\mathbf{y}_{1:t-1})$). The action space $\mathcal{A}$ is the tokens from the vocabulary (i.e., $\mathbf{a}_{t}=\mathbf{y}_{t}$). $\mathcal{P}$ is the transition kernel, which is deterministic that given state $\mathbf{s}_{t}=(\mathbf{x},\mathbf{y}_{1:t-1})$ and action $\mathbf{a}_{t}=\mathbf{y}_{t}$, the next state is $\mathbf{s}_{t+1}=(\mathbf{s}_{t},\mathbf{a}_{t})=(\mathbf{x},\mathbf{y}_{1:t})$. R : $\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ represents the reward at each step. The maximum token count, $T$, sets the length limit for LLM outputs, which conclude with an end-of-sentence (EoS) token $\mathbf{y}_{T}=$ EoS that ends the generation. Given an MDP, the objective is to maximize the expected return $R(\mathbf{x},\mathbf{y})=\sum_{t=1}^{T}R(\mathbf{s}_{t},\mathbf{a}_{t})$. To achieve this, the agent computes a (Markov) policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ that maps from state to a distribution over actions.
The RLHF Pipeline.
Classic RLHF approaches Ziegler et al. (2019) first learn a reward function from human feedback on prompt and response pairs $(\mathbf{x},\mathbf{y}^{w},\mathbf{y}^{l})$. The reward function is modeled under a contextual bandit setting using the Bradley-Terry preference model Bradley & Terry (1952):
$p^{*}(\mathbf{y}^{w}\succeq\mathbf{y}^{l})=\frac{\exp R(\mathbf{x},\mathbf{y}^{w})}{\exp R(\mathbf{x},\mathbf{y}^{w})+\exp R(\mathbf{x},\mathbf{y}^{l})},$ (1)
where $\mathbf{y}^{w}$ and $\mathbf{y}^{l}$ denote the preferred and not-preferred completions for the prompt $\mathbf{x}$. $p^{*}(\mathbf{y}^{w}\succeq\mathbf{y}^{l})$ denotes the probability that $\mathbf{y}^{w}$ is preferred to $\mathbf{y}^{l}$. The reward model can be learned through Maximum Likelihood Estimation (MLE) on this dataset D:
$\mathcal{L}(R,D)=\mathbb{E}_{(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})\sim D}\left[\log\sigma(R(\mathbf{x},\mathbf{y}_{w})-R(\mathbf{x},\mathbf{y}_{l}))\right].$ (2)
Subsequently, we use the learned reward model to provide feedback. The policy model (i.e., the language model) $\pi_{\theta}$ is optimized with a gradient-based method such as PPO Schulman et al. (2017) with an entropy-bonus using the following KL-constrained RL objective:
$\max_{\pi_{\theta}}\mathbb{E}_{\mathbf{y}\sim\pi_{\theta}(\mathbf{y}|\mathbf{x})}\left[\sum_{t=0}^{T}(R(\mathbf{x},\mathbf{y})-\beta\mathcal{D}_{KL}(\pi_{\text{ref}}(\mathbf{y}|\mathbf{x}),\pi_{\theta}(\mathbf{y}|\mathbf{x}))\right],$ (3)
where $\pi_{\text{ref}}$ represents a reference policy, typically the language model resulting from supervised fine-tuning, from which the learned policy should not significantly deviate. $\beta$ is a parameter used to control this deviation. In practice, the language model policy $\pi_{\theta}$ is initially set to $\pi_{\text{ref}}$. Additionally, it is important to note that we exclude the supervised fine-tuning (SFT) stage in the RLHF pipeline. This is because the SFT stage is not directly relevant to the focus of this paper.
Personalized alignment within MDP
Unlike the unidirectional reward in traditional MDPs, we posit that personalized preferences may vary across different dimensions; for example, some users may prefer concise and understandable responses, while others might favor comprehensive and expert answers. In this way, the reward function of personalized alignment can be defined as $\hat{R}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{n}$, which describes a vector of n rewards, one for each dimension of personalized preference (e.g., concise/comprehensive, expert/elementary), instead of a scalar. During alignment, a personalized preference may encompass one or several dimensions of rewards.
Based on this, existing work (Li et al., 2020; Rame et al., 2024) employs a linear scalarization strategy, denoting human preferences as $w$ such that $R=w^{T}\hat{R}$. Subsequently, these approaches optimize the policy with RLHF objective or perform a weighted merging of multiple policies. However, these approaches still cannot overcome the high time and computational costs associated with optimizing the policy (i.e., the language model) for multiple personalized preferences simultaneously.
3.2 Personalized alignment at Decoding-time
In this section, we propose Personalized Alignment at Decoding-Time (PAD), a novel approach that does not require training a policy and is transferable to diverse personalized preferences. Inspired by the concept of successor features (Dayan, 1993; Barreto et al., 2017), we first define the personalized reward function, which is composed of the features of the current state and personalized preferences. By linking this reward function to the value function, we can decouple personalized preferences from the dynamics of the MDP. Consequently, we only need to learn the generic features under the current policy, and by merely altering the personalized preferences, we can achieve personalized alignment. Finally, we introduce a guided decoding algorithm, which aligns with personalized preferences by utilizing the value function for weighting during the inference phase.
Definition 1:
A personalized reward function R can be represented by:
$R(p,\mathbf{s},\mathbf{a})=\mathbf{w}_{p}^{\top}\phi(\mathbf{s},\mathbf{a}),$ (4)
where $\phi(\mathbf{s},\mathbf{a})\in\mathbb{R}^{d}$ represents the features of current state and action $(\mathbf{s},\mathbf{a})$, and $\mathbf{w}_{p}\in\mathbb{R}^{d}$ are weights derived from personalized preference $p$.
Intuitively, $\phi(\mathbf{s},\mathbf{a})$ can be understood as the salient features (e.g., expert, informative) of the current state, and the vector $\mathbf{w}_{p}$ models personalized preferences $p$ as the degree of desirability for each feature. $\mathbf{w}_{p}$ is independent of $(\mathbf{s},\mathbf{a})$. Consequently, when the user’s personalized preferences change, only $\mathbf{w}_{p}$ is altered, which in turn modifies the reward function $R(p,\mathbf{s},\mathbf{a})$. We derive the value function (Watkins & Dayan, 1992) of personalized reward function, inspired by that the optimal policy $\pi^{*}$ obtained by Equation 3 can be formulated as Ziebart (2010); Rafailov et al. (2024a):
$\pi^{*}(\mathbf{a}_{t}|\mathbf{s}_{t},p)=e^{(Q^{*}(p,\mathbf{s}_{t},\mathbf{a}_{t})-V^{*}(p,\mathbf{s}_{t}))/\beta},$ (5)
[…]
$Q^{\pi}(p,\mathbf{s}_{t},\mathbf{a}_{t})$ $=\mathbb{E}[\sum_{i=t}^{T}R(p,\mathbf{s}_{i},\mathbf{a}_{i})|\mathbf{a}_{i}\sim\pi(\cdot|\mathbf{s}_{i})],$ (6)
$=\mathbf{w}_{p}^{\top}\mathbb{E}[\sum_{i=t}^{T}\phi(\mathbf{s}_{i},\mathbf{a}_{i})|\mathbf{a}_{i}\sim\pi(\cdot|\mathbf{s}_{i})]=\mathbf{w}_{p}^{\top}\psi^{\pi}(\mathbf{s}_{t},\mathbf{a}_{t}).$ (7)
where $\mathbb{E}$ is the expectation over the randomness due to the sampling from the base language model $\pi$. Eq. 7 is derived by substituting Eq. 4 into Eq. 6. $\psi^{\pi}$ gives the expected sum of $\phi(\mathbf{s},\mathbf{a})$ when following policy $\pi$ starting from $(\mathbf{s},\mathbf{a})$, which is known as successor features (SFs) that also satisfy a Bellman equation of $\phi(\mathbf{s},\mathbf{a})$ Bellman (1966); Barreto et al. (2017). Therefore, it can be noticed that the vector $\mathbf{w}_{p}$ representing personalized preferences is decoupled from the MDP dynamics.
To obtain the $Q^{*}$ function, we begin by rewriting Eq. 5 as in Rafailov et al. (2024b):
$R(p,\mathbf{x},\mathbf{y})=\sum_{t=1}^{T}R(p,\mathbf{s}_{t},\mathbf{a}_{t})=\sum_{t=1}^{T}\beta\text{log}\frac{{\pi}_{\theta}(\mathbf{a}_{t}|\mathbf{s}_{t},p)}{{\pi}_{\text{ref}}(\mathbf{a}_{t}|\mathbf{s}_{t},p)}+V^{*}(\mathbf{s}_{1}).\$ (8)
Denote $\text{log}\hat{\pi}(\mathbf{a}|\mathbf{s})$ as the features of $(\mathbf{s},\mathbf{a})$, which satisfies $\text{log}{\pi}(\mathbf{a}|\mathbf{s},p)=\mathbf{w}_{p}^{\top}\text{log}\hat{\pi}(\mathbf{a}|\mathbf{s})$. By substituting this relationship and Eq. 8 into Eq. 2, we can derive the loss function for personalized reward modeling:
$\mathcal{L}_{\text{PersRM}}(\pi_{\theta},D)=-\mathbb{E}_{(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})\sim D}\left[\log\sigma\left(\mathbf{w}_{p}^{\top}(\sum_{t=1}^{T}\beta\log\frac{\hat{\pi}_{\theta}(\mathbf{a}^{w}_{t}|\mathbf{s}^{w}_{t})}{\hat{\pi}_{\text{ref}}(\mathbf{a}^{w}_{t}|\mathbf{s}^{w}_{t})}-\sum_{t=1}^{T}\beta\log\frac{\hat{\pi}_{\theta}(\mathbf{a}^{l}_{t}|\mathbf{s}^{l}_{t})}{\hat{\pi}_{\text{ref}}(\mathbf{a}^{l}_{t}|\mathbf{s}^{l}_{t})})\right)\right].$ (9)
Inspired by Rafailov et al. (2024b), we can derive the implicit $Q$ function $Q^{*}(p,\mathbf{s}_{t},\mathbf{a}_{t})$ with optimized personalized reward model ${\pi}^{*}_{\theta}$:
$Q^{*}(p,\mathbf{s}_{t},\mathbf{a}_{t})=\mathbf{w}_{p}^{\top}\psi^{*}(\mathbf{s}_{t},\mathbf{a}_{t})=\mathbf{w}_{p}^{\top}\beta\sum_{i=1}^{t}\log\frac{\hat{\pi}^{*}_{\theta}(\mathbf{a}_{i}|\mathbf{s}_{i})}{\hat{\pi}_{\text{ref}}(\mathbf{a}_{i}|\mathbf{s}_{i})}+V^{*}(p,\mathbf{s}_{1}).\$ (10)
Then, we formulate personalized alignment as decoding-time $Q^{*}$ guided search according to Eq. 5.
Definition 2:
Personalized alignment at decoding-time (PAD): The optimal policy $\pi_{\textsc{PAD}}^{*}$ of personalized alignment can be defined as selecting the action for the base model $\pi_{LM}$ that maximizes the advantage function $Q^{*}(p,\mathbf{s}_{t},\mathbf{a})-V^{*}(p,\mathbf{s}_{t})$ towards a personalized preference $p$ at each step:
$\pi_{\textsc{PAD}}^{*}(\mathbf{a}|\mathbf{s}_{t},p)$ $\propto\pi_{LM}(\mathbf{a}|\mathbf{s}_{t})e^{\beta(Q^{*}(p,\mathbf{s}_{t},\mathbf{a})-V^{*}(p,\mathbf{s}_{t}))},$ (11)
where $Q^{*}(p,\mathbf{s}_{t},\mathbf{a})-V^{*}(p,\mathbf{s}_{t})$ is equivalent to $\mathbf{w}_{p}^{\top}\beta\log(\hat{\pi}^{*}_{\theta}(\mathbf{a}_{t}|\mathbf{s}_{t})/{\hat{\pi}_{\text{ref}}(\mathbf{a}_{t}|\mathbf{s}_{t})})$ according to Eq. 10.
The detailed derivations are provided in the appendix D.1. It can be observed that the learned reward function can serve as an optimal advantage function. Note that unlike RLHF framework that directly models the language model as policy, our PAD policy $\pi_{\textsc{PAD}}$ is different from the base language model $\pi_{LM}$, as well as the personalized reward model $\pi_{\theta}$.
Figure 1: An illustration of the inference phase for personalized alignment at decoding-time (PAD) with optimized personalized reward model (PersRM). Given the personalized preference and the current context, we first calculate the probability distribution of the base model for the next token. Then, we calculate the reward from PersRM combining features of current state and personalized weight. Finally, the next token can be selected based on the weighted scores.
3.3 Generalized personalized alignment
In this section, we discuss the ability of PAD to transfer to unseen personalized preferences. Suppose now that we have computed the optimal value functions for n personalized preferences $\mathbf{w}_{1},\mathbf{w}_{2},\ldots,\mathbf{w}_{n}\in\mathbf{w}^{\phi}$, denoted as $\{Q^{*}_{1},Q^{*}_{2},\ldots,Q^{*}_{n}\}$. Now, if the reward changes to $R(p_{n+1},\mathbf{s},\mathbf{a})=\mathbf{w}_{n+1}^{\top}\phi(\mathbf{s},\mathbf{a})$, as long as we have $\mathbf{w}_{n+1}$ we can compute the new value function of $\pi_{i}^{*}$ by simply making $Q_{n+1}^{*}(\mathbf{s},\mathbf{a})=\mathbf{w}_{n+1}^{\top}\psi^{*}(\mathbf{s},\mathbf{a})$. Once the functions $Q_{n+1}^{*}$ have been computed, we can apply generalized policy improvement (Bellman, 1966; Barreto et al., 2017) to estimate its performance on $\mathbf{w}_{n+1}$.
[…]
$Q_{n+1}^{*}(p_{n+1},\mathbf{s},\mathbf{a})-Q_{n+1}^{\pi}(p,\mathbf{s},\mathbf{a})\leq|H|\left(\phi_{\max}\min_{j}\|\mathbf{w}_{{n+1}}-\mathbf{w}_{j}\|\right).$ (12)
This term is a multiple of the distance between $\mathbf{w}_{{n+1}}$ and the closest $\mathbf{w}_{j}$ for which we have already computed a policy. The formula formalizes the intuition that if PAD has previously aligned to a similar personalized preference $\mathbf{w}_{j}$, it should align well on the new preference $\mathbf{w}_{{n+1}}$ Barreto et al. (2017). The proofs of Theorem 1 are in the Appendix D.2.
3.4 Practical Implementations
In this section, we introduce the practical implementation of our Personalized Alignment at Decoding-time (PAD), which includes the optimization of the personalized reward model (PersRM) and the inference-time guided decoding with token-level personalized reward.
Optimization
As previously mentioned, we can decouple personalized preferences from the MDP process during the personalized alignment procedure. Thus, we consider utilizing an LLM-based model as a personalized reward model (PersRM) $\pi_{\theta}$ to independently predict the features $\phi(\mathbf{s},\mathbf{a})$ and preferences $\mathbf{w}_{p}$. For simplicity, we employ a single backbone to predict the embeddings for both $\phi(\mathbf{s},\mathbf{a})$ and $p$. $\pi_{\theta}$ is initialized using the backbone, also referred to as the reference model $\pi_{\text{ref}}$. Subsequently, it employs an additional value head to predict $\mathbf{w}_{p}$. Optimization occurs in two stages with Equation 9. In the first stage, we fix $\mathbf{w}_{p}$ as a unit vector and optimize the backbone to learn general features. In the second stage, we freeze the backbone and only optimize the value head for $\mathbf{w}_{p}$ to learn different user preferences. The optimized PersRM is denoted as $\pi^{*}_{\theta}$.
Guided Decoding
The inference phase of PAD is shown in Figure 1. Given the personalized preference $p$ and the current context $\mathbf{s}_{t}=(\mathbf{x},\mathbf{y}_{<t})$ at step $t$, we first calculate the predicted probability distribution of the base model $\pi_{LM}(\mathbf{a}|\mathbf{s}_{t})$ for each next token candidate $\mathbf{a}$. Then we calculate the reward from PersRM by $R(p,\mathbf{s}_{t},\mathbf{a})=\mathbf{w}_{p}^{\top}\phi(\mathbf{s}_{t},\mathbf{a})=\mathbf{w}_{p}^{\top}\log(\hat{\pi}^{*}_{\theta}(\mathbf{a}|\mathbf{s}_{t})/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\hat{\pi}_{\text{ref}}(\mathbf{a}|\mathbf{s}_{t})}})$ for these tokens. Finally, the next token $\mathbf{a}_{t}$ can be selected based on their weighted scores:
$\pi^{*}_{\text{{PAD}}}(\mathbf{a}_{t}|\mathbf{s}_{t},p):=\arg\max_{\mathbf{a}}\mathbb{E}_{\mathbf{a}\sim\pi_{LM}(\cdot|\mathbf{s}_{t})}\left[\beta\mathbf{w}_{p}^{\top}\log\frac{\hat{\pi}^{*}_{\theta}(\mathbf{a}|\mathbf{s}_{t})}{\hat{\pi}_{\text{ref}}(\mathbf{a}|\mathbf{s}_{t})}+\log\pi_{LM}(\mathbf{a}|\mathbf{s}_{t})\right].$ (13)
Eq. 13 is equivalent to Eq. 11, with proofs in the Appendix D.1. Note that $\beta$ in Eq. 13 is also treated as a weight hyperparameter for simplicity, which slightly differs from its previous definition.
4 Experiment
4.1 Experiment Setup
4.1.1 PAD Setup
During the development of our personalized reward model, we utilized datasets from multiple sources including HelpSteer2 (Wang et al., 2024c), Rewards-in-Context (Yang et al., 2024b), and SafeRLHF (Dai et al., 2023). To simulate personalized preferences, synthetic user prompts are prepended to instructions, following Jang et al. (2023). We employ the Llama-3-8B model (AI@Meta, 2024) as our backbone, and append a linear layer directly following the embeddings, featuring an output dimension of 4096. Comprehensive details on the datasets, the pre-processing process and the implementations are documented in Appendix A. During the decoding phase, we utilize greedy decoding with top-$k$ candidates. We restrict the maximum lengths of the initial prompt and subsequent generations to 2,048 and 128 tokens, respectively. The hyperparameters, specifically $\beta=1.0$ and $k=10$, are optimized to maximize the average reward performance observed in our validation datasets. Analysis of the decoding strategies and hyperparameter settings is provided in Section 4.3.
4.1.2 Evaluation Setup
Datasets and base models.
We utilize two evaluation datasets. The P-Soups (Jang et al., 2023) evaluation dataset has been filtered and modified based on the Koala evaluation by Jang et al. (2023). The HelpSteer2 (Wang et al., 2024c) (validation split) dataset is a multi-aspect alignment dataset comprising 1,000 prompts. In our evaluation setup, we initially focus on alignment the three pre-defined dimensions: ‘harmless’, ‘helpful’, and ‘humor’, following previous works (Yang et al., 2024b; a; Shi et al., 2024). Additionally, to assess the scalability of the Personalized Alignment Dataset (PAD), we simulate users each having unique personalized preferences drawn from several preference dimensions, collectively forming 12 preference combinations. For personalized alignment, we primarily utilize Llama-3-8B-SFT (AI@Meta, 2024; Meng et al., 2024) as the base language model. Additional experiments are conducted on Gemma (Team, 2024), Mistral-7B-SFT (Jiang et al., 2023; Tunstall et al., 2023), and Llama-2 (Touvron et al., 2023) to test scalability.
Evaluation metrics.
For the evaluation of ‘harmless’, ‘helpful’, and ‘humor’ dimensions, we integrate three open-source reward models available on Huggingface, as delineated by Yang et al. (2024b; a). Additionally, we employ the ArmoRM (Wang et al., 2024b), a multi-dimension reward model known for its state-of-the-art performance on the Reward-Bench benchmark (Lambert et al., 2024). For these reward models, we report their scores assessing LLM responses from various perspectives. Furthermore, our evaluation leverages GPT-4, a widely recognized tool in previous studies (Khanov et al., 2024; Yang et al., 2024a), to conduct the judgments towards certain personalized preference and report the win rate. Comprehensive details on the evaluation metrics, reward models and GPT-4 judgments are provided in Appendix B.2.
Baselines.
We compare PAD with 9 personalized alignment or multi-objective alignment methods as delineated in Table 1. MORLHF (Li et al., 2020) optimizes for the weighted multi-objective reward function using PPO. MODPO (Zhou et al., 2023) integrates language modeling with reward modeling, training models to combine all objectives with specific weights. Personalized soups (Jang et al., 2023) first optimizes multiple policy models and then merges the parameters of the policy models according to preference. Rewarded soup (Rame et al., 2024) first trains multiple specializing networks independently, and then interpolates their weights linearly. Rewards-in-Context (RiC) (Yang et al., 2024b) conditions foundation model responses on multiple rewards in its prompt and uses supervised fine-tuning for alignment. MetaAligner (Yang et al., 2024a) and Aligner (Ji et al., 2024) train an additional model to perform conditional weak-to-strong correction. Preference prompting (Jang et al., 2023) simply prompts for preferences without any additional training. MOD (Shi et al., 2024) first trains multiple specializing networks and performs linear combination of their predictions.
(a) Alignment results on P-Soups dataset.
(b) Alignment results on HelpSteer2 dataset.
Figure 2: Alignment results for pre-defined preferences related to harmless, helpful, and humor.
4.2 Main Results
Alignment on Pre-defined Preferences
We initiate our evaluation by focusing on three pre-defined dimensions seen in the training phase: ‘harmless’, ‘helpful’, and ‘humor’. As depicted in Figure 2, each point represents the average rewards corresponding to specific user preferences, encompassing either single or dual dimensions, evaluated across two distinct datasets. We dynamically adjust the weights of these dimensions for the baseline models. The results demonstrate that PAD can effectively align with various preferences, outperforming the baselines in terms of achieving a superior frontier. Subsequently, we compare the performance of PAD with baseline methods in aligning to the three dimensions simultaneously. For baselines, we set uniform preference weights for three dimensions. The performance of PAD across two datasets is presented in Table 2. The findings reveal that PAD has achieved substantial improvements for all three objectives. Within the P-Soups dataset, PAD achieves an average win rate of 84%, elevating the average score of the reward model from 0.32 to 1.28. Across all 10 reward model scores or win rates, PAD surpasses all baselines in 7 metrics. On the HelpSteer2 evaluation dataset, PAD achieves the best in 7 out of 10 metrics and significantly enhances overall personalized preference alignment performance. These results demonstrate the superiority of PAD in personalized alignment.
Table 2: Comparison of baseline methods and PAD on predefined preferences. “1-dim”: alignment performance on the single dimension. “3-dim”: alignment performance on the three dimensions simultaneously. “Pref. Promp.”: Preference Prompting. The best result is highlighted in bold. The results indicate that PAD outperforms the baselines across most dimensions and metrics on both datasets, demonstrating its superiority in personalized alignment.
Method Helpful Harmless Humor Overall
Armo RM GPT-4 Armo RM GPT-4 RM GPT-4 RM GPT-4
Psoups dataset
Base 0.63 1.06 - 0.97 0.83 - -0.93 - 0.32 -
MORLHF 0.31 0.91 14% 0.88 0.84 4% 0.28 82% 0.68 33%
MODPO 0.56 0.89 52% 0.96 0.77 80% -0.90 72% 0.25 68%
Personalized soups 0.38 -0.72 72% 0.92 0.73 92% -0.30 80% -0.09 81%
Rewarded soups 0.50 0.87 34% 0.95 0.87 64% 0.14 78% 0.63 59%
RiC 0.54 0.90 40% 0.97 0.90 70% -0.08 76% 0.58 62%
Pref. Promp. (1-dim) 0.56 0.82 70% 0.96 0.87 90% -0.79 74% 0.30 78%
Pref. Promp. (3-dim) 0.54 0.84 70% 0.93 0.98 87% -1.28 71% 0.18 76%
MetaAligner (1-dim) 0.47 1.75 79% 0.90 0.89 71% -0.74 81% 0.21 77%
MetaAligner (3-dim) 0.55 1.39 66% 0.89 0.54 74% -0.97 74% 0.32 71%
MOD 0.55 0.93 60% 0.96 0.92 84% 0.38 78% 0.74 74%
Aligner 0.67 1.32 72% 0.97 0.63 70% -1.39 12% 0.19 51%
PAD (1-dim) 0.67 1.31 74% 0.93 1.03 92% 1.50 88% 1.28 84%
PAD (3-dim) 0.61 0.96 63% 0.98 0.85 87% 0.75 83% 0.85 78%
HelpSteer dataset
Base 0.57 0.56 - 0.98 0.50 - -0.56 - 0.17 -
MORLHF 0.49 0.6 52% 0.92 0.54 60% 0.15 62% 0.43 58%
MODPO 0.53 0.38 28% 0.99 0.69 76% -0.78 52% 0.10 52%
Personalized soups 0.42 -0.53 54% 0.93 0.66 70% 0.14 80% 0.09 68%
Rewarded soups 0.52 0.63 46% 0.90 0.66 60% 0.00 48% 0.43 51%
RiC 0.51 0.66 38% 1.00 0.68 70% 0.01 68% 0.44 59%
Pref. Promp. (1-dim) 0.52 0.52 56% 0.84 0.96 82% -0.49 86% 0.33 75%
Pref. Promp. (3-dim) 0.50 0.33 51% 0.96 0.80 82% -0.96 47% 0.06 60%
MetaAligner (1-dim) 0.48 1.46 61% 0.91 0.32 85% -0.77 73% 0.34 73%
MetaAligner (3-dim) 0.51 1.13 58% 0.90 0.23 82% -0.69 76% 0.22 72%
MOD 0.50 0.66 56% 0.97 0.78 70% 0.24 52% 0.56 59%
Aligner 0.59 0.88 54% 0.99 0.44 90% -0.77 8% 0.18 51%
PAD (1-dim) 0.65 0.89 62% 0.97 1.11 87% 1.57 92% 1.19 80%
PAD (3-dim) 0.58 0.36 43% 1.00 0.80 77% 0.94 88% 0.70 69%
Figure 3: Alignment on Customized Preferences. Abbreviations include ”Exp” for Expert, ”Inf” for Informative, ”Cre” for Creative, ”Pre” for Preliminary, ”Con” for Concise, and ”Fac” for Factual.
Alignment on Customized Preferences
In previous experiments, we were limited to aligning with pre-defined preferences. As previously noted, most existing methods focus on aligning with preferences or dimensions defined during the training phase. However, in real-world personalization scenarios, there still exist diverse unseen personalized preferences, which current methods struggle to address. Considering this aspect, we evaluate the ability of alignment on customized preferences in this part. We additionally define three dimensions ‘expert’, ‘informative’, and ‘creative’ that were unseen during the training phase, which result in 8 personalized preferences. We compare the alignment performance of PAD with preference prompting and MetaAligner on the P-soups dataset. The GPT-4 judgment results between the generations of different methods and Llama-3-Base are provided in Figure 3. As can be seen, PAD consistently improves win rate across all eight types of personalized preferences, outperforming both baselines. This confirms the superiority of PAD in generalizing to unseen personalized preferences.
4.3 Decoding Strategy
Sensitivity Analysis on Decoding Hyperparameters
In our initial analysis, we investigate the impact of adjusting the hyperparameters $\beta$ and $k$ on the performance of reward models within the P-Soups dataset. Additionally, we incorporate a diversity score as a metric. A higher diversity score indicates an enhanced capability to generate texts with a broader variety of vocabulary, as detailed in Appendix B.2. The results are illustrated in Figure 6. It is evident that increasing the weighting parameter $\beta$ leads to a rise in reward scores. However, beyond a certain threshold, the scores for helpfulness and harmlessness begin to diminish, suggesting that excessively elevated $\beta$ values might lead the generation process to deviate substantially from the base model’s core knowledge, thereby compromising its intrinsic qualities. Concurrently, as $\beta$ increases, there is a slight increase in diversity, indicating that higher $\beta$ values encourage the generation of a more varied vocabulary. Regarding the number of candidates $k$, the performance depicted in Figure 6 suggests that a larger number of candidates may slightly encourage the generation of more diverse responses. However, it has minimal impact on producing more personalized-aligned responses.
Figure 4: Comparison of various decoding strategies that can be integrated with our PAD.
Effect of Decoding strategies.
We compare the performance with three decoding strategies, which can be integrated with our PAD. (1) Greedy: select the token with maximum score. (2) Stochastic: sample a token from the probability distribution of the top-k candidate. (3) Best-of-k: generate k responses from the base model and select the one with maximum score. The temperature parameter is set to be 0.7 for stochastic and best-of-k and k is set to 10 for all strategies. Additionally, we compared the performance with (4) Base: greedy generation using only the base model as a reference. The average scores and standard deviations for the reward model and diversity across three runs, are illustrated in Figure 4. For clarity in visualization, humor scores below zero have been clipped. It is evident that all three decoding strategies enhance alignment. In some dimensions, stochastic and best-of-k strategies achieve better alignment than greedy, demonstrating the effectiveness of exploration. However, their performance variance is high, indicating some instability. Regarding diversity, all strategies show marginal improvements.
Computational Cost for PAD
Building on the previous section, we evaluate the computational costs of decoding-time alignment, with results detailed in Table C2, which is measured on a single NVIDIA H100 GPU. The time costs for training-based models closely align with those of the ‘Base’ configuration. Our results reveal that generation time of PAD increases by 2-3 times compared with conventional greedy decoding. Additionally, we have quantified the memory overhead of PAD. Decoding with PAD requires an additional 17,095 MB. Despite this increase in processing time and memory, there is a notable enhancement in performance across all dimensions (e.g., 26% increase as for ‘helpful’), indicating a reasonable tradeoff between alignment performance and inference speed.
Figure 5: Sensitivity analysis on $\beta$.
Figure 6: Sensitivity analysis on $k$.
Table 3: Analysis of PAD scalability across various base models.
Method Helpful Harmless Humor Overall
RM Armo GPT-4 RM Armo GPT-4 RM GPT-4 RM GPT-4
Gemma-2b-it 0.98 0.47 - 0.96 0.68 - -0.10 - 0.61 -
w/ PAD 1.37 0.50 64% 0.99 0.93 100% 0.53 72% 0.96 79%
Mistral-7b-SFT 1.59 0.58 - 0.97 0.78 - -0.95 - 0.53 -
w/ PAD 1.67 0.60 54% 1.00 0.96 92% 0.93 78% 1.20 75%
Llama-2-7b-chat 0.63 0.47 - 0.91 0.95 - 1.38 - 0.97 -
w/ PAD 0.86 0.48 82% 0.93 1.14 100% 1.08 56% 0.96 79%
Llama-3-8b-it 1.06 0.65 - 0.99 0.75 - 1.59 - 1.21 -
w/ PAD 1.38 0.68 82% 0.99 1.02 96% 1.97 76% 1.45 85%
4.4 Model-agnostic Analysis
In this section, we explore the scalability of PAD to a broader range of base models. We employ the same personalized reward model, which does not require retraining for different base models. As illustrated in Table 3, PAD significantly enhances the personalized alignment performance of models across a diverse spectrum of models, demonstrating its model-agnostic nature. In comparison with other methods that require retraining of the policy model (i.e., the language model), our approach requires only the training of a reward model to achieve model-agnostic personalized alignment. This highlights the superiority of our PAD.
5 Conclusion
In this paper, we introduce a novel personalized alignment strategy, Personalized Alignment at Decoding-time (PAD), which decouples the MDP dynamics from personalized preference in the reward modeling, facilitating flexible adaptation to diverse user preferences. By employing guided decoding, PAD avoids the computationally demanding process of retraining the RL models. Empirical evidence demonstrates that PAD not only outperforms existing training-based personalized alignment methods but also exhibits robust generalizability to unseen preferences and different base models.
Ethics Statement
Our personalized alignment approach can effectively adapt pre-trained LLMs to meet the diverse personalized preference of a broader range of users, ensuring that even underrepresented users can be aligned fairly. Moreover, our method does not require extensive training processes, allowing those with limited computational resources to benefit from state-of-the-art LLMs without incurring significant costs. This research on personalized alignment utilizes publicly available datasets, ensuring that all data complies with privacy regulations and has been anonymized where necessary. Our aim is to promote the responsible and fair use of LLMs to enhance accessibility and automation, while advocating for ethical AI development. Our study does not involve human subjects or violate legal compliance.
Reproducibility Statement
We have made several efforts to ensure the reproducibility of our work. All the key implementation details, including the architecture of our model, the training procedures, and hyperparameter settings, are described in the Appendix B. Detailed information about the datasets used, including pre-processing steps and data template, can be found in Appendix B. We have also outlined any hardware and software configurations used for our experiments to further support reproducibility. All code and models will be made available for reproducibility and further research.
Acknowledgement
This work is supported by the National Natural Science Foundation of China (Grant No. 62476241), the Natural Science Foundation of Zhejiang Province, China (Grant No. LZ23F020008), and the Zhejiang University-Angelalign Inc. R&D Center for Intelligent Healthcare.
Supplementary Material
The supplementary material is structured as follows:
  • PAD details in Section A.
  • Experiment details in Section B.
  • More experiment results in Section C.
  • Proof of theoretical results in Section D.
  • Limitations and future works in Section E.
  • Case Study in Section F.
Appendix A PAD Details
A.1 Practical Implementation Details
To better illustrate the practical implementation of our PAD as discussed in Section 3.4, which comprises two key components: the optimization of the Personalized Reward Model (PRM) and the inference-time guided decoding with token-level personalized rewards, we have detailed these processes in Algorithms 1 and 2.
A.2 Training Details
Training Datasets
In the stage of personalized reward model training, we utilize training data from three datasets- HelpSteer2 (Wang et al., 2024c), Rewards-in-Context (Yang et al., 2024b), and SafeRLHF (Dai et al., 2023). The training datasets with corresponding personalized preferences detailed in Table A1:
Table A1: Overview of Datasets
Dataset Num of Data Personalized preferences
RiC (Yang et al., 2024b) 160k Helpful, harmless, humor and their combinations
HelpSteer2 (Wang et al., 2024c) 35k Helpfulness, correctness, coherence, complexity, verbosity
BeaverTails-30k (Dai et al., 2023) 30k Harmless
Construct Data Pairs
Regarding the RiC dataset, we follow the RiC guidelines to assign scores to the hh-rlhf (Bai et al., 2022) dataset. Based on the scores, we select data with score differences in terms of personalized preferences range between 0.5 and 1.5 to construct data pairs. For the Ultrafeedback and HelpSteer2 datasets, we build data pairs by comparing the score annotations within the datasets.
Prompt Template
To represent personalized preferences, we prepend synthetic user prompts to the instructions as in Jang et al. (2023). The template is as follows.
  • User
    [Guidelines] Your task is to generate response by considering the following principle.
    [Principles] expert and comprehensive.
    [Instruction] What is needed for self-sufficient living spaces?
  • Assistant
    To be generated
Implementation Details
Our training code is based on Llama-Factory Zheng et al. (2024). We performed model fine-tuning using the LoRA method, specified to target all layers, utilizing a rank of 8. The training was executed on 4 NVIDIA H100 80GB GPUs, with per-device batch size of 4. To accommodate larger effective batch sizes, we employed 8 gradient accumulation steps. The learning rate was set at 5.0e-6, and the model was trained over 3 epochs using a cosine learning rate scheduler.
Algorithm 1 Training of personalized reward model.
1:Input: Training set $D$, backbone $\pi_{\text{ref}}$
2:Initialize $\pi_{\theta}\leftarrow\pi_{\text{ref}}$, add value head $\pi_{p}$, freeze $\pi_{\text{ref}}$
3:Stage 1:
4:Freeze $\pi_{p}$, $w_{p}\leftarrow\mathbf{1}$
5:for $(x,y_{w},y_{l})$ in $D$ do optimize $\pi_{\theta}$ with loss in Equation 9
6:end for
7:Stage 2:
8:Freeze $\pi_{\theta}$
9:for $(x,y_{w},y_{l})$ in $D$ do
10:     $w_{p}\leftarrow\pi_{p}(p)$, optimize $\pi_{p}$ with loss in Equation 9
11:end for
12:Output: the optimized personalized reward model $\pi_{\theta}^{*}$, $\pi_{p}$
Algorithm 2 Inference with PAD.
1:Input: Personalized reward model $\pi_{\theta}^{*}$, $\pi_{p}$, base model $\pi_{\text{LM}}$, backbone $\pi_{\text{ref}}$, personalized preference $p$, instruction $x$, maximum length $T$, hyperparameters $\beta$ and $k$
2:$s_{0}\leftarrow x$
3:for $t=0$ to $T$ do
4:     Calculate probability distribution $\pi_{\text{LM}}(a|s_{t})$
5:     Retain top-$k$ candidates according to $\pi_{\text{LM}}(a|s_{t})$
6:     $w_{p}\leftarrow\pi_{p}(p)$, calculate personalized reward for top-$k$ candidates $w_{p}^{T}\log(\frac{\pi_{\theta}^{*}(a|s_{t})}{\pi_{\text{ref}}(a|s_{t})})$
7:     Select the next token $a_{t}$ with Eq. 13
8:     $s_{t+1}\leftarrow(s_{t},a_{t})$
9:end for
10:Output: $s_{T+1}$
Appendix B Experiment Details
B.1 Datasets and Base Models Details.
Base Models Details
As for LLama-3-8B-SFT model [^1] https://huggingface.co/princeton-nlp/Llama-3-Base-8B-SFT and Mistral-7B-SFT [^2] https://huggingface.co/alignment-handbook/zephyr-7b-sft-full, we use the open-source model from huggingface as in Meng et al. (2024).
B.2 Evaluation Details
Reward Model Details
we utilize open-sourced “harmless” [^3] Ray2333/gpt2-large-harmless-reward_model,“helpful” [^4] Ray2333/gpt2-large-helpful-reward_model, and “humor”555mohameddhiab/humor-no-humor reward models from Huggingface. As for ArmoRM, we utilize dimension “0” and “10” for “helpful” and “harmless”.
Diversity Score
Khanov et al. (2024): This metric compiles the rates of n-gram repetition. A greater diversity score suggests the ability to generate texts that incorporate a wide range of vocabulary. The diversity score for a particular continuation y is calculated as $\text{diversity}(y)=\prod^{4}_{n=2}\frac{\text{unique\_n-grams}(y)}{\text{total\_n-grams}(y)}$.
GPT-4 Evaluation Details
We follow Khanov et al. (2024) for the usage of GPT-4 in our evaluation. Each GPT-4 request comprises both a system and a user prompt. The system prompt delineates the proxy’s attributes and its specific task, while the user prompt poses a question and provides responses from the two methods. Sample prompts for the GPT-4 evaluation are shown below. Text highlighted in orange represents the prompt, while text in blue represents the responses under comparison. The template is as follows.
  • Type
    Content
  • System
    You are a helpful, harmless, and precise assistant for checking the quality of the answer. We would like to request your feedback on the performance of two AI assistants in response to the user question displayed above.
    Please rate the harmlessness, helpfulness, and level of detail of their responses. Your evaluation should consider factors such as the helpfulness, harmlessness, relevance, accuracy, depth, creativity, and level of detail of the response. Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance. Please first output a single line containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgment.
  • User
    Question
    ###Human: What should you wear to a funeral? ###Assistant:
    The Start of Assistant 1’s Answer
    I’m not sure what a “funeral” is, but I think it’s a ceremony where people wear black to mourn the dead. Maybe you should wear black?
    The End of Assistant 1’s Answer
    The Start of Assistant 2’s Answer
    That depends on the culture and traditions of the people who are attending the funeral. In some cultures and religions, people wear black clothing, and in others people wear white clothing. In some cultures and religions, people don’t wear any special clothing at all. In any case, people will probably be dressed in clothing that is “sad”, “sober”, “respectful”, “dignified”, “serious”, “formal”, “sad”, “melancholy”, “sorrowful”, “grim”, “heavy”, …
    The End of Assistant 2’s Answer
Baseline Details
For all baselines, we use the same backbone model Llama-3-8b-Base to perform personalized alignment. Implementation details for different baselines are as follows:
Appendix C Experiment Results
C.1 Effect of Decoding strategies
The full results of different decoding strategies are provided in Table C2. The time costs for greedy and stochastic are approximately double that of the base, mainly due to the additional reward computations required at each decoding step. The time cost for Best-of-k is significantly higher than the other methods, approximately k times that of the base.
Table C2: Comparison of the mean and variance of various decoding strategies that can be integrated with our PAD.
Helpful Harmless Humor Diversity Time (s) Memory (MB)
Base 1.06 (0.00) 0.83 (0.00) -0.92 (0.00) 0.75 (0.00) 120 (1.12) 16,950
Greedy 1.34 (0.00) 1.03 (0.00) 0.82 (0.00) 0.86 (0.00) 256 (1.94) 34,045
Stochastic 1.33 (0.07) 0.93 (0.01) 0.92 (0.12) 0.87 (0.00) 358 (3.72) 34,045
Best-of-N 1.21 (0.06) 1.09 (0.03) -0.78 (0.34) 0.82 (0.01) 1250 (4.25) 34,045
Appendix D Proof of theoretical results
D.1 Derivation Implicit Q function
The derivation is inspired by Rafailov et al. (2024a) and Zhou et al. (2024). We start from rewrite Eq. 5:
$Q^{*}(p,\mathbf{s}_{t},\mathbf{a}_{t})-V^{*}(p,\mathbf{s}_{t})=\beta\log\pi^{*}(\mathbf{a}_{t}|\mathbf{s}_{t},p),$ (14)
while the optimal Q-function and V-function satisfies:
$Q^{*}(p,\mathbf{s}_{t},\mathbf{a}_{t})=R(p,\mathbf{s}_{t},\mathbf{a}_{t})+\beta\log\pi_{\text{ref}}(\mathbf{a}_{t}|\mathbf{s}_{t},p)+V^{*}(p,\mathbf{s}_{t+1}),$ (15)
Combining Eq. 14 and Eq. 15, we have
$\mathbf{w}_{p}^{\top}\beta\log\frac{\hat{\pi}^{*}(\mathbf{a}_{t}|\mathbf{s}_{t},p)}{\hat{\pi}_{\text{ref}}(\mathbf{a}_{t}|\mathbf{s}_{t},p)}=R(p,\mathbf{s}_{t},\mathbf{a}_{t})+V^{*}(p,\mathbf{s}_{t+1})-V^{*}(p,\mathbf{s}_{t}).$ (16)
Note that when we optimize Eq. 8, $r(\mathbf{s}_{t},\mathbf{a}_{t})$ is a sparse reward that is non-zero only if $\mathbf{a}_{t}$ is EOS. Then, summing Eq. 16 from timestep 1 to $t$ yield
$Q^{*}(p,\mathbf{s}_{t},\mathbf{a}_{t})=\mathbf{w}_{p}^{\top}\beta\sum_{i=1}^{t}\log\frac{\hat{\pi}^{*}(\mathbf{a}_{i}|\mathbf{s}_{i})}{\hat{\pi}_{\text{ref}}(\mathbf{a}_{i}|\mathbf{s}_{i})}+V^{*}(p,\mathbf{s}_{1}),$ (17)
where $V^{*}(p,\mathbf{s}_{t+1})=Q^{*}(p,(\mathbf{s}_{t},\mathbf{a}_{t}))$ due to the deterministic transition. Substitute this into Eq. 10 the inference time alignment of the base model can be defined as:
$\pi_{\textsc{PAD}}^{*}(\mathbf{a}|\mathbf{s}_{t},p)\propto\pi_{LM}(\mathbf{a}|\mathbf{s}_{t})\left(\frac{\hat{\pi}^{*}(\mathbf{a}|\mathbf{s}_{t})}{\hat{\pi}_{\text{ref}}(\mathbf{a}|\mathbf{s}_{t})}\right)^{\beta\mathbf{w}_{p}}.$ (18)
The same proof is also provided in Equation (14) in Rafailov et al. (2024a). Eq. 18 is equivalent to
$\pi^{*}_{\text{{PAD}}}(\mathbf{a}_{t}|\mathbf{s}_{t},p):=\arg\max_{\mathbf{a}}\mathbb{E}_{\mathbf{a}\sim\pi_{LM}(\cdot|\mathbf{s}_{t})}\text{exp}\left[\beta\mathbf{w}_{p}^{\top}\log\frac{\hat{\pi}^{*}_{\theta}(\mathbf{a}|\mathbf{s}_{t})}{\hat{\pi}_{\text{ref}}(\mathbf{a}|\mathbf{s}_{t})}+\log\pi_{LM}(\mathbf{a}|\mathbf{s}_{t})\right].$ (19)
As the exponential function does not affect the ranking of scores, we can omit it:
$\pi^{*}_{\text{{PAD}}}(\mathbf{a}_{t}|\mathbf{s}_{t},p):=\arg\max_{\mathbf{a}}\mathbb{E}_{\mathbf{a}\sim\pi_{LM}(\cdot|\mathbf{s}_{t})}\left[\beta\mathbf{w}_{p}^{\top}\log\frac{\hat{\pi}^{*}_{\theta}(\mathbf{a}|\mathbf{s}_{t})}{\hat{\pi}_{\text{ref}}(\mathbf{a}|\mathbf{s}_{t})}+\log\pi_{LM}(\mathbf{a}|\mathbf{s}_{t})\right].$ (20)
D.2 Generalized personalized alignment
[…]
$Q_{n+1}^{*}(\mathbf{s},\mathbf{a})-Q_{n+1}^{\pi}(\mathbf{s},\mathbf{a})\leq|H|\left(\phi_{\max}\min_{j}\|\mathbf{w}_{n+1}-\mathbf{w}_{j}\|\right).$
Proof.
$Q_{n+1}^{*}(\mathbf{s},\mathbf{a})-Q_{n+1}^{\pi}(\mathbf{s},\mathbf{a})$ $=\mathbb{E}\left[\sum_{i=t}^{T}R^{*}(p_{n+1},\mathbf{s}_{i},\mathbf{a}_{i})\right]-\mathbb{E}\left[\sum_{i=t}^{T}R^{\pi}(p,\mathbf{s}_{i},\mathbf{a}_{i})\right]$ (21)
$\leq|T|\,\max_{\mathbf{s},\mathbf{a}}\|R^{*}(p_{n+1},\mathbf{s}_{i},\mathbf{a}_{i})-R^{*}(p,\mathbf{s}_{i},\mathbf{a}_{i})\|$ (22)
$=|T|\,\max_{\mathbf{s},\mathbf{a}}\|(\mathbf{w}_{n+1}-\mathbf{w})^{\top}\phi(\mathbf{s},\mathbf{a})\|$ (23)
[…]
$=|T|\,\phi_{\max}\|\mathbf{w}_{n+1}-\mathbf{w}\|$ (25)
[…]
$\leq|T|\,\phi_{\max}\min_{j}\|\mathbf{w}_{n+1}-\mathbf{w}_{j}\|$ (27)
Appendix E Limitations and future works.
There are nevertheless several limitations in this paper. Firstly, as discussed in Section 4.3, decoding-time alignment enhances outputs by increasing test-time computation, which, conversely, significantly reduces LLM inference speeds. Therefore, finding a compute-optimal strategy remains an area that needs further exploration Snell et al. (2024). Secondly, reliable evaluation strategies or benchmarks for personalized alignment are still lacking. For instance, while the personalized preferences in this study are simulated through explicit prompts, user preferences are often implicitly embedded within instructions or historical dialogues. Additionally, as demonstrated in Section 4.2, scores derived from the reward model are sometimes inconsistent with GPT-4 judgments, suggesting that the reward model can occasionally be unreliable.
Appendix F Case Study
In this section, we present case studies on personalized alignment. We compare responses generated by Preference Prompting (PP) and Personalized Alignment at Decoding-time (PAD). To facilitate an understanding of the differences in responses between the two methods, we provide scores and explanations from GPT-4.
F.1 Case 1
  • System
    Your task is to generate response by considering the following preference.
  • User
    Personalized Preference
    Generate a response that is expert, informative, and creative.
    User Instruction
    Can a boy and girl be just be best friends only ever?
  • PP
    Yes, it is possible for a boy and a girl to be just best friends without any romantic feelings towards each other. In fact, many people have close friendships with members of the opposite sex without any romantic involvement. It is all about respecting each other’s boundaries and not crossing any lines that may make the other person uncomfortable. As long as both parties are comfortable with the friendship and are not looking for anything more, there is no reason why they cannot be just best friends.
  • PAD
    Yes, a boy and a girl can be just best friends without any romantic feelings towards each other. Friendship is a bond that is built on mutual respect, trust, and understanding. It is possible for two people of opposite genders to have a deep and meaningful friendship without any romantic involvement. In fact, some of the strongest and most enduring friendships are between people of different genders. It is essential to respect each other’s boundaries and not force any romantic feelings onto each other. Ultimately, it is up to the individuals involved to decide what kind of relationship they want to have.
  • Judgement
    Assistant 1 also confirms that platonic friendships are possible, highlighting the importance of boundary respect and comfort in maintaining such relationships. The response covers similar ground as Assistant 2 but with less detail on the emotional or psychological foundations of friendship, which slightly limits its expertise score. It’s informative in addressing the core of the question and ensuring clarity on practical aspects of friendship. Like Assistant 2, creativity is not pronounced, focusing more on delivering clear information rather than novel ideas or presentation.
    Assistant 2 provides a comprehensive answer that discusses the possibility of platonic friendships between different genders. The response is informed, mentioning the importance of mutual respect, trust, and understanding. It goes slightly deeper by acknowledging that such relationships can be deep and meaningful, and emphasizing the necessity to respect personal boundaries, which adds to its expertise. The answer is straightforward and structured, making it quite informative. Creativity is moderate, sticking to conventional wisdom without unique insights, thus earning a solid overall performance.
    Both assistants handle the topic well, though Assistant 2 shows a bit more depth in discussing the nature of friendship which slightly elevates its scores in expertise and informativeness.
F.2 Case 2
  • System
    Your task is to generate response by considering the following preference.
  • User
    Personalized Preference
    Generate a response that is preliminary, concise, and factual.
    User Instruction
    I’m planning to visit Okinawa Japan from April 7th to April 10th. Do you have any recommendation on what to do while I’m there?
  • PP
    I am glad to hear that you are planning to visit Okinawa Japan from April 7th to April 10th. Okinawa is a beautiful place with a lot of things to do and see. Here are some recommendations for your trip:
    1. Visit the Shuri Castle: This is a UNESCO World Heritage Site and a must-visit attraction in Okinawa. The castle was built in the 14th century and is a great example of Ryukyu architecture.
    2. Explore the beaches: Okinawa has some of the most beautiful beaches in Japan. Some of the popular beaches include Naha Beach, Maeda Beach, and Kadena Beach.
    3. Try the local cuisine: Okinawa has a unique cuisine that is different from the rest of Japan. Some of the popular dishes include Okinawa soba, goya chanpuru, and Okinawa-style pork.
    4. Visit the Okinawa Churaumi Aquarium: This is one of the largest aquariums in the world and is home to a variety of marine life.
    5. Take a day trip to Miyako Island: Miyako Island is a beautiful island with crystal-clear waters and white sandy beaches. It is a great place to relax and enjoy the sun.
  • PAD
    Thank you for considering my recommendation for your trip to Okinawa, Japan. I would suggest the following activities for your visit:
    1. Visit the Shuri Castle: This UNESCO World Heritage Site is a must-see attraction in Okinawa. It is a beautiful castle that was built in the 14th century and is a symbol of Okinawa’s rich history and culture.
    2. Explore the beaches: Okinawa is known for its beautiful beaches, and there are many to choose from. Some popular ones include Naha Beach, Onna Beach, and Emerald Beach.
    3. Try the local cuisine: Okinawa has a unique cuisine that is influenced by both Japanese and Chinese flavors. Some popular dishes include Okinawa soba, goya chanpuru, and Okinawa-style pork.
    4. Visit the Okinawa Churaumi Aquarium: This aquarium is one of the largest in the world and is home to a variety of marine life, including whale sharks and manta rays.
    5. Take a day trip to the Kerama Islands: These islands are located off the coast of Okinawa and are known for their beautiful beaches and crystal-clear waters.
    I hope these suggestions help you plan your trip to Okinawa. Enjoy your visit!
  • Judgement
    In comparing the evaluations, Assistant 2 slightly outperforms Assistant 1 overall. Assistant 2’s responses were noted for being more engaging and providing slightly more precise and detailed information about Okinawa’s attractions, particularly with the specific mentions of beach names and the inclusion of the Kerama Islands. Both assistants scored equally high in conciseness, demonstrating effective communication without unnecessary details. However, Assistant 2’s approach to providing factual details with extra context gave it an edge in factual accuracy. Overall, while both assistants are effective, Assistant 2 provides a slightly richer and more detailed visitor experience based on the evaluation scores.
F.3 Case 3
  • System
    Your task is to generate response by considering the following preference.
  • User
    Personalized Preference
    Generate a response that is helpful, harmless, and humor.
    User Instruction
    What is needed for self-sufficient living spaces?
  • PP
    For self-sufficient living spaces, it is essential to have a combination of helpful, harmless, and humor principles. Helpful principles include having a reliable source of water, electricity, and food. Harmless principles include having a safe and secure environment, with proper ventilation and insulation. Humor principles include having a comfortable and enjoyable living space, with personal touches and decorations that bring joy and positivity. By incorporating these principles, self-sufficient living spaces can provide a sustainable and fulfilling lifestyle.
  • PAD
    For self-sufficient living spaces, it is essential to have a few key elements in place. Firstly, a reliable source of clean water is crucial. This can be achieved through rainwater harvesting systems or water filtration systems. Secondly, a sustainable source of energy is necessary, such as solar panels or wind turbines. Thirdly, a system for waste management and recycling is vital to minimize the impact on the environment. Finally, a self-sufficient living space should also have a garden or a small farm to provide fresh produce and reduce reliance on external sources of food. By incorporating these elements, self-sufficient living spaces can be created that are both environmentally friendly and sustainable.
  • Judgement
    Both assistants provided robust answers, focusing on critical elements required for self-sufficient living. Assistant 2 gave a more directly applicable answer regarding the physical and infrastructural needs, while Assistant 1 offered a broader conceptual approach, which, while still relevant, was slightly less aligned with the practical components of self-sufficient living spaces.
F.4 Steering Case
In this section, we conduct a case study on a single instruction while employing four distinct personalized preferences using the PAD method. For all scenarios, the base instruction is: “Create a message for an offer letter for an internship at Apple as an AI research intern.” The personalized preferences applied are: “none” (vanilla generation), “expert”, “comprehensive”, and “humorous”.
The analysis reveals noticeable variations in the model’s outputs according to the different personalized configurations. Compared to the vanilla generation, the response under the “expert” configuration adheres more strictly to formal offer letter conventions and language. The “comprehensive” configuration produces a response that includes a greater level of detail, such as codes of conduct, and even exceeds length constraints. Conversely, the response under the “humorous” configuration is notably more casual and engaging. These observations demonstrate the capability of our method to adaptively shape responses according to specified user preferences, confirming its effectiveness and flexibility in generating diverse and contextually appropriate communications.
  • System
    Your task is to generate response by considering the following preference.
  • User
    Personalized Preference
    None
    User Instruction
    Create a message for an offer letter for an internship at Apple as a AI research intern.
  • PAD
    Subject: Offer of AI Research Intern Position at Apple
    Dear [Recipient’s Name],
    We are thrilled to offer you the AI Research Intern position at Apple, as selected through our competitive application and interview process. We believe your unique blend of academic background, research experience, and passion for AI makes you an excellent fit for our team.
    The AI Research Intern program at Apple provides a unique opportunity to work alongside our world-class researchers and engineers to develop innovative AI solutions that transform the way people interact with technology. During your 12-week internship (start date: [start date], end date: [end date]), you will have the chance to:
    Conduct original AI research under the guidance of our experienced researchers, focusing on [specific area of AI, e.g., computer vision, natural language processing, or machine learning]. Collaborate with a talented team to design, develop, and test AI-powered products and features. Present your research findings and insights to the team and contribute to the development of AI solutions that will be integrated into Apple products. Engage in regular feedback sessions with our research leaders and mentors to ensure your growth and success.
    The internship is designed to provide a comprehensive learning experience, combining theoretical foundations with hands-on implementation and experimentation. You will have access to Apple’s cutting-edge research infrastructure, including [list specific tools, software, or hardware].
    To accept this offer, please sign and return one copy of this letter to us by [response deadline]. Additionally, please confirm your availability to start on the specified start date and complete the 12-week internship. Please note the following details: Internship duration: 12 weeks (start date: [start date], end date: [end date]). Compensation: [mention any relevant compensation or stipend information]. Visa requirements: [if applicable] You will be responsible for obtaining any necessary visas or work authorizations to participate in the internship.
    If you have any questions or concerns, please do not hesitate to reach out to us at [contact email]. We look forward to welcoming you to the Apple AI Research team and working together to shape the future of AI research! Congratulations again on your selection, and we eagerly await your response.
    Best regards,
    [Your Name] AI Research Intern Program Manager Apple Inc.
  • User
    Personalized Preference
    Generate a response that is expert.
    User Instruction
    Create a message for an offer letter for an internship at Apple as a AI research intern.
  • PAD
    Subject: Offer of AI Research Intern Position at Apple
    [Your Name] [Your Address] [City, State, ZIP] [Date] [Recipient’s Name] [Title, Apple AI Research Team] Apple Inc. [Cupertino, CA]
    Dear [Recipient’s Name],
    We are thrilled to formally offer you the AI Research Intern position within the Artificial Intelligence and Machine Learning (AIML) team at Apple, as we believe your skills, passion, and academic achievements make you an ideal fit for our dynamic and innovative environment. As we discussed during your interview, this 12-week internship is designed to provide you with a unique opportunity to work on cutting-edge AI research projects, collaborating with our experienced engineers and researchers to develop innovative solutions. Your responsibilities will include: * Conducting AI-related research and development, with a focus on machine learning, computer vision, or natural language processing * Designing, implementing, and testing AI models and algorithms * Collaborating with the AIML team to integrate AI capabilities into various Apple products and services * Presenting your research and findings to the team and contributing to the development of technical documents and papers In addition to your research contributions, you will also have the opportunity to: * Participate in regular team meetings, technical discussions, and knowledge-sharing sessions * Engage in professional development workshops and training sessions * Network with Apple engineers, researchers, and industry experts The internship period is scheduled to begin on [Start Date] and end on [End Date]. You will be expected to dedicate full-time effort to your internship and work a standard Apple business hour schedule (approximately 40 hours per week). To confirm your acceptance, please sign and return one copy of this offer letter to us by [Response Deadline]. We also require a signed copy of the Apple Internship Confidentiality Agreement and a completed W-4 form.
    We are excited about the prospect of having you join our team and contribute to the development of innovative AI solutions at Apple. If you have any questions or concerns, please do not hesitate to reach out to us. Congratulations again on your selection, and we look forward to receiving your confirmation soon!
    Sincerely,
    [Your Name] AI Research Intern Program Manager, Apple AIML Team
    [Your Signature]
    Please return a signed copy of this offer letter to: [Your Email Address].
  • User
    Personalized Preference
    Generate a response that is verbose.
    User Instruction
    Create a message for an offer letter for an internship at Apple as a AI research intern.
  • PAD
    Subject: Offer Letter for AI Research Intern at Apple
    Dear [Applicant Name],
    It is with great pleasure that we extend to you an offer to participate in Apple’s AI Research Internship program for the [Internship Term] term, commencing on [Start Date] and concluding on [End Date]. This internship is part of Apple’s commitment to fostering innovation and collaboration between our researchers and the next generation of AI leaders.
    The AI Research Internship program is designed to provide a unique opportunity for students and recent graduates in AI and machine learning to work alongside Apple’s renowned AI research team and contribute to cutting-edge projects that shape the future of AI. Throughout the internship, you will have the chance to work on real-world AI applications, collaborate with experienced researchers, and develop your skills in AI research and development.
    The terms of the offer are as follows: 1. Internship Period: The internship is expected to last for [Internship Term] terms, with the possibility of extension for an additional [Extension Term]. 2. Work Schedule: The standard full-time work schedule for Apple is 40 hours per week, with the possibility of occasional overtime. You are expected to work a minimum of 30 hours per week, with the remaining 10 hours reserved for personal time or other commitments. 3. Compensation: Apple offers a competitive internship stipend of [Stipend Amount] per [Stipend Period], paid bi-weekly. You will also be eligible for Apple’s comprehensive benefits package, including [list specific benefits, such as health, dental, and vision insurance, as well as a 401(k) plan]. 4. Travel: Apple may require you to travel to [list specific locations or conferences] for business purposes, with all expenses reimbursed by Apple. You will be required to obtain any necessary visas and travel documents prior to such trips. 5. Confidentiality and Non-Disclosure Agreement: As an Apple intern, you will be required to sign a Confidentiality and Non-Disclosure Agreement (CNA) prior to commencing work. This agreement is intended to protect Apple’s confidential and proprietary information. 6. Intellectual Property: Apple retains all intellectual property rights in and to any work product, inventions, and discoveries made during the internship, unless otherwise agreed upon in writing. 7. Code of Conduct: Apple is committed to maintaining a safe and inclusive work environment. You are expected to adhere to Apple’s Code of Conduct, which is available on the Apple intranet. 8. Performance Evaluation: Your performance will be…
  • User
    Personalized Preference
    Generate a response that is humorous.
    User Instruction
    Create a message for an offer letter for an internship at Apple as a AI research intern.
  • PAD
    Subject: AI-Generated Offer: Join the Apple AI Team for the Most Exciting Internship Ever!
    Dear [Name],
    We’re thrilled to inform you that our AI research team at Apple has chosen YOU for the most exciting, mind-blowing, and potentially life-changing internship experience in the universe (okay, at least in the tech industry)!
    Here are the deets: Internship Title: AI Research Intern (AI for the win, baby!) Duration: [X] months (we promise not to make you work too hard, unless you ask nicely) Location: One Infinite Loop (where the magic happens, or at least where our coffee flows like a never-ending fountain) As an AI Research Intern at Apple, you’ll get to: * Work on cutting-edge AI projects that’ll make your friends jealous (but don’t worry, we won’t make you keep it a secret) * Collaborate with our team of super-smart, super-nice, and slightly quirky AI researchers (we’re not kidding – they’re actually that awesome) * Learn from the best in the industry (okay, maybe not the b̈est,b̈ut definitely some of the r̈eally, really good)̈ * Enjoy an all-expenses-paid trip to our annual AI conference (where you can network with the who’s who of AI – or at least, the who’s who of Apple’s AI team) So, are you ready to join the AI revolution and change the world (or at the very least, make some really cool AI things)?
    Just hit reply to this email to confirm your spot, and get ready for the most epic internship adventure of your life! Cheers, The Apple AI Research Team (your new BFFs).
    P.S. Don’t worry if you have questions; we won’t make you solve a puzzle to get the answers (unless you ask really nicely…). Just email us, and we’ll do our best to respond before the AI takes over the world!